Two point charges q and ?q are separated by a distance L. What is the magnitude of the force on a third point charge q placed at the midpoint between them?

(a) 8kq2/L2

(b) 2kq2/L2

(c) 0

(d) 4kq2/L2

(e) kq2/L2

any explanation is helpful!
Answer #1


let q1 = q

q2 = -q


q3 = q ( at the middle)


lert r1 = r2 = L/2


force exerted by 1 on 3

F13 = k*q1*q/r1^2 = k*q^2*4/L^2 this force is repulsive and
is direted towards q2

fore exerted by 2 on 3


F23 = k*q2*q/r2^2 = k*q^2*4/L^2 this force is attrative and
is directed towards q2

Fnet = F13 + F23 = k*q^2*4/L^2 + k*q^2*4/L^2

Fnet = 8kq2/L2

option A